# Model-Based Learning for DOA Estimation with One-Bit Single-Snapshot Sparse Arrays

### v1

@@ III-B One-Bit Off-Grid DOA Estimation via SBRI Framework
$p\left(\overline{\mathbf{y}}|\hat{\mathbf{x}};\bm{\beta}\right)=\!\prod_{m=1}^{M}$ $\Phi\left(\Re(\overline{y}_{m})\Re\left(\bm{c}_{m}^{\mathrm{T}}(\bm{\beta})\hat{\mathbf{x}}\right)\right)$
×$\times\Phi\left(\Im(\overline{y}_{m})\Im\left(\bm{c}_{m}^{\mathrm{T}}(\bm{\beta})\hat{\mathbf{x}}\right)\right).$ (20)
The off-grid gaps $\bm{\beta}$ follow a uniform distribution, $p\left(\bm{\beta}\right)\sim U\left(-\frac{r}{2},\frac{r}{2}\right)$[50], independent of $\hat{\mathbf{x}}$, where $r$ denotes the grid interval size. Using Bayes’ rule, the MAP estimator for the off-grid model is given by:
@@
Substituting (20) and (10) into (21), we obtain the following cost function to be minimized:
$\mathcal{L}_{\mathrm{offgrid}}$ $=\sum_{m=1}^{M}\Bigg{\{}-\mathrm{ln}\,\Phi\left(\Re\left(\overline{y}_{m}\right)\Re(\bm{c}_{m}^{\mathrm{T}}(\bm{\beta})\hat{\mathbf{x}})\right)$
$\ -\mathrm{ln}\,\Phi\left(\Im\left(\overline{y}_{m}\right)\Im(\bm{c}_{m}^{\mathrm{T}}(\bm{\beta})\hat{\mathbf{x}})\right)\!+\!\sum_{i=1}^{N}\frac{\gamma\left|x_{i}\right|^{\alpha}}{\alpha}\!+\!\textrm{const}\Bigg{\}}.$ (22)
@@
where $\mathbf{w}^{k}=[w_{1}^{k},\cdots,w_{M}^{k}]^{\mathrm{T}}$, $w_{m}^{k}=\left[\Re(\overline{y}_{m})\Re(\tilde{w}_{m}^{k})\right]+j\left[\Im(\overline{y}_{m})\Im(\tilde{w}_{m}^{k})\right]$, $\tilde{w}_{m}^{k}=e_{m}^{k}-\mathrm{I}^{\prime}(e_{m}^{k})$, $e_{m}^{k}=\Re(\overline{y}_{m})\Re(\bm{c}_{m}^{\mathrm{T}}(\bm{\beta}^{k})\hat{\mathbf{x}}^{k})+j\Im(\overline{y}_{m})\Im(\bm{c}_{m}^{\mathrm{T}}(\bm{\beta}^{k})\hat{\mathbf{x}}^{k})$. By employing an alternating iterative approach, we update the spectral coefficients $\hat{\mathbf{x}}$ and the grid gaps $\bm{\beta}$ as follows:
@@
×$\times\Re\left(\mathrm{diag}\left((\hat{\mathbf{x}}^{k+1})^{*}\right)\mathcal{B}^{\mathrm{H}}\left[\mathbf{w}^{k}-\mathcal{A}\hat{\mathbf{x}}^{k+1}\right]\right).$ (25)
The derivations for (24) and (25) are provided in the Appendix. The proposed SBRI for the one-bit off-grid model is summarized in Algorithm 2.
@@
3:Initialize $\hat{\mathbf{x}}^{0}=\frac{\mathcal{A}^{H}\overline{\mathbf{y}}}{\|\mathcal{A}^{H}\overline{\mathbf{y}}\|_{2}}$, $\bm{\beta}^{0}=\bm{0}$
4:for $k\leftarrow 1$ to $T_{\max}$ do
5:     Update $\hat{\mathbf{x}}^{k}$ using the formula in (24)
6:     Update $\bm{\beta}^{k}$ using the formula in (25)
7:     Update $\gamma^{k}$ using the formula in (19)
8:     if ($\frac{\|\hat{\mathbf{x}}^{(k+1)}-\hat{\mathbf{x}}^{(k)}\|_{2}^{2}}{\|\hat{\mathbf{x}}^{(k)}\|_{2}^{2}}\leq\epsilon_{0}$ and $\frac{\|\bm{\beta}^{k+1}-\bm{\beta}^{k}\|_{2}^{2}}{\|\bm{\beta}^{k}\|_{2}^{2}}\leq\epsilon_{0}$) or $k\geq T_{\max}$ then
9:         break
10:     else
11:         $\hat{\mathbf{x}}^{k+1}\leftarrow\hat{\mathbf{x}}^{k}$
12:         $\bm{\beta}^{k+1}\leftarrow\bm{\beta}^{k}$
13:         $\gamma^{k+1}\leftarrow\gamma^{k}$
14:     end if
15:end for
16:return $\hat{\mathbf{x}}^{\star}\leftarrow\hat{\mathbf{x}}^{k}$, $\bm{\beta}^{\star}\leftarrow\bm{\beta}^{k}$
@@ III-C One-Bit DOA Estimation via Augmented Framework: SBRI-X
$p\left(\tilde{\mathbf{x}}\mid\overline{\mathbf{y}},\bm{\epsilon}\right)\propto p\left(\overline{\mathbf{y}}\mid\tilde{\mathbf{x}},\bm{\epsilon}\right)p\left(\tilde{\mathbf{x}}\right).$ (31)
@@
$\left(\tilde{\mathbf{x}}^{\star},\bm{\epsilon}^{\star}\right)=\mathrm{arg}\max_{\tilde{\mathbf{x}},\bm{\epsilon}}p\left(\overline{\mathbf{y}}\mid\tilde{\mathbf{x}},\bm{\epsilon}\right)p\left(\tilde{\mathbf{x}}\right).$ (32)
Following the same approach as in Section III-A, taking the negative logarithm of (32) yields the following non-convex objective function:
@@
$=$ $\sum_{i=1}^{N}\ln{\left(1+\mathrm{a}\exp\left(-\mathrm{b}\overline{y}_{i}\left[\mathbf{a}_{i}^{\mathrm{T}}\tilde{\mathbf{x}}+\epsilon_{i}\right]\right)\right)}$
@@
III-C1 Majorizing Function for $\mathcal{L}^{\prime}\left(\tilde{\mathbf{x}},\bm{\epsilon}\right)$
@@
By substituting (36) into (33) and applying smooth approximation on the second term in (33), we can obtain the majorizing function ¯$\bar{\mathcal{L}}_{ongrid}^{\diamond}\left(\tilde{\mathbf{x}},\bm{\epsilon}\right)$ of $\mathcal{L}_{ongrid}^{\diamond}\left(\tilde{\mathbf{x}},\bm{\epsilon}\right)$:
@@ III-C2 Updating Formula
2:Output: $\hat{\mathbf{x}}^{\star}$
4:for $k\leftarrow 1$ to $T_{\max}$ do
5:     Update $\tilde{\mathbf{x}}^{k}$ using the formula in (41)
6:     Update $\bm{\epsilon}^{k}$ using the formula in (42)
7:     Update $\gamma^{k}$ using the formula in (19)
8:     if ($\frac{\|\hat{\mathbf{x}}^{(k+1)}\!\!-\hat{\mathbf{x}}^{(k)}\|_{2}^{2}}{\|\hat{\mathbf{x}}^{(k)}\|_{2}^{2}}\!\leq\!\epsilon_{0}$ and $\frac{\|\bm{\epsilon}^{k+1}\!\!-\bm{\epsilon}^{k}\|_{2}^{2}}{\|\bm{\epsilon}^{k}\|_{2}^{2}}\!\leq\!\epsilon_{0})$ or $k\!\geq\!T_{\mathrm{max}}$ then
9:         break
10:     else
11:         $\hat{\mathbf{x}}^{k+1}\leftarrow\hat{\mathbf{x}}^{k}$
12:         $\bm{\epsilon}^{k+1}\leftarrow\bm{\epsilon}^{k}$
13:         $\gamma^{k+1}\leftarrow\gamma^{k}$
14:     end if
15:end for
16:return $\hat{\mathbf{x}}^{\star}\leftarrow\hat{\mathbf{x}}^{k}$
@@
$\tilde{\mathbf{x}}^{k+1}=$ $\left[\mathbf{C}^{\mathrm{H}}(\bm{\beta}^{k})\mathbf{C}(\bm{\beta}^{k})+\frac{\gamma(\mathrm{a}+1)^{2}}{\mathrm{a}\mathrm{b}^{2}}\bm{\Lambda}^{-1}(\tilde{\mathbf{x}}^{k})\right]^{-1}$
@@
$g^{\dagger}(\tilde{\mathbf{x}}^{k},\bm{\epsilon}^{k})=$ $\frac{(\mathrm{a}+1)^{2}\Re({\overline{\mathbf{y}}})}{\mathrm{b}\exp\left(\mathrm{b}\Re({\overline{\mathbf{y}}})\odot\Re\left(\mathbf{C}(\bm{\beta}^{k})\tilde{\mathbf{x}}^{k}+\bm{\epsilon}^{k}\right)\right)+\mathrm{a}\mathrm{b}}$
@@
2:Output: $\hat{\mathbf{x}}^{\star}$
3:Initialize $\tilde{\mathbf{x}}^{0}=\frac{\mathcal{A}^{H}\overline{\mathbf{y}}}{\|\mathcal{A}^{H}\overline{\mathbf{y}}\|_{2}}$, $\bm{\beta}^{0}=\mathbf{0}$
4:for $k\leftarrow 1$ to $T_{\max}$ do
5:     Update $\tilde{\mathbf{x}}^{k}$ using the formula in (43)
6:     Update $\bm{\epsilon}^{k}$ using the formula in (45)
7:     Update $\bm{\beta}^{k}$ using the formulas in (46) and (47)
8:     Update $\lambda^{k}$ using the formula in (19)
9:     if ($\frac{\|\hat{\mathbf{x}}^{(k+1)}\!\!-\hat{\mathbf{x}}^{(k)}\|_{2}^{2}}{\|\hat{\mathbf{x}}^{(k)}\|_{2}^{2}}\!\leq\!\epsilon_{0}$ and $\frac{\|\bm{\epsilon}^{k+1}\!\!-\bm{\epsilon}^{k}\|_{2}^{2}}{\|\bm{\epsilon}^{k}\|_{2}^{2}}\!\leq\!\epsilon_{0}$) or $k\!\geq\!T_{\mathrm{max}}$ then
10:         break
11:     else
12:         $\tilde{\mathbf{x}}^{k+1}\leftarrow\tilde{\mathbf{x}}^{k}$
13:         $\bm{\epsilon}^{k+1}\leftarrow\bm{\epsilon}^{k}$
14:         $\bm{\beta}^{k+1}\leftarrow\bm{\beta}^{k}$
15:         $\gamma^{k+1}\leftarrow\gamma^{k}$
16:     end if
17:end for
18:return $\tilde{\mathbf{x}}^{\star}\leftarrow\hat{\mathbf{x}}^{k}$

### v2

@@ III-B One-Bit Off-Grid DOA Estimation via SBRI Framework
[…]
The off-grid gaps $\bm{\beta}$ follow a uniform distribution, $p\left(\bm{\beta}\right)\sim U\left(-\frac{r}{2},\frac{r}{2}\right)$[55], independent of $\hat{\mathbf{x}}$, where $r$ denotes the size of the grid interval. We adopt a uniform prior for the off-grid gaps $\bm{\beta}$ for three reasons. First, with no deterministic knowledge of the true DOAs, the uniform distribution is the non-informative (maximum-entropy) choice, assigning equal likelihood across each grid cell. Second, this matches the assumptions of leading off-grid sparse Bayesian learning methods, e.g., OGSBI [55], off-grid SBL-RVM [56], and OGIR [37], thus ensuring a fair comparison in Section V. Finally, the uniform prior admits simple closed-form MAP updates and directly yields the efficient update rules in Equations (25)–(27) and Equations (50)–(53), whereas alternatives (e.g., Gaussian or Beta priors) would introduce extra hyperparameters and necessitate numerical integration. Using Bayes’ rule, the MAP estimator for the off-grid model is given by:
@@
Substituting (21) and (10) into (22), we obtain the following cost function to minimize:
[…]
$\ -\mathrm{ln}\,\Phi\left(\Im\left(\overline{y}_{m}\right)\Im(\bm{c}_{m}^{\mathrm{T}}(\bm{\beta})\hat{\mathbf{x}})\right)\left.\!+\!\sum_{i=1}^{N}\frac{\gamma\left|x_{i}\right|^{\alpha}}{\alpha}\!+\!\textrm{const}\right\}.$ (23)
@@
where $\mathbf{w}^{k}=[w_{1}^{k},\cdots,w_{M}^{k}]^{\mathrm{T}}$, $w_{m}^{k}=\left[\Re(\overline{y}_{m})\Re(\tilde{w}_{m}^{k})\right]+j\left[\Im(\overline{y}_{m})\Im(\tilde{w}_{m}^{k})\right]$, $\tilde{w}_{m}^{k}=e_{m}^{k}-\mathrm{I}^{\prime}(e_{m}^{k})$, $e_{m}^{k}=\Re(\overline{y}_{m})\Re(\bm{c}_{m}^{\mathrm{T}}(\bm{\beta}^{k})\hat{\mathbf{x}}^{k})+j\Im(\overline{y}_{m})\Im(\bm{c}_{m}^{\mathrm{T}}(\bm{\beta}^{k})\hat{\mathbf{x}}^{k})$. By employing an alternating iterative approach, we first compute $\mathbf{w}^{k}$ and then update the spectral coefficients $\hat{\mathbf{x}}$ and the grid gaps $\bm{\beta}$ as follows:
$\mathbf{w}^{k}$ $=[\Re(\overline{\mathbf{y}})\odot\Re(\mathbf{E}^{k}-\mathrm{I}^{\prime}(\mathbf{E}^{k}))]+j[\Im(\overline{\mathbf{y}})\odot\Im(\mathbf{E}^{k}-\mathrm{I}^{\prime}(\mathbf{E}^{k}))],$ (25)
@@
$\times\Re\left(\mathrm{diag}\left((\hat{\mathbf{x}}^{k+1})^{*}\right)\mathcal{B}^{\mathrm{H}}\left[\mathbf{w}^{k}-\mathcal{A}\hat{\mathbf{x}}^{k+1}\right]\right),$ (27)
where $\mathbf{E}^{k}=[\Re(\overline{\mathbf{y}})\odot\Re(\mathbf{C}(\bm{\beta}^{k})\hat{\mathbf{x}}^{k})]+j[\Im(\overline{\mathbf{y}})\odot\Im(\mathbf{C}(\bm{\beta}^{k})\hat{\mathbf{x}}^{k})].$
The derivations for (26) and (27) are as follows. First, we compute the first-order derivative of the objective function in Equation (24) with respect to $\hat{\mathbf{x}}$ as:
$\frac{\partial{\mathcal{L}}\left(\hat{\mathbf{x}},\bm{\beta}^{k}\right)}{\partial\hat{\mathbf{x}}}$= $\mathbf{C}^{\mathrm{H}}(\bm{\beta}^{k})\left[\mathbf{C}(\bm{\beta}^{k})\hat{\mathbf{x}}-\mathbf{w}^{k}\right]+\gamma^{k}\bm{\Lambda}(\hat{\mathbf{x}})\hat{\mathbf{x}}$.
Setting the derivative to zero, we obtain:
[…]
The solution of (29) yields Equation (26). With the estimate $\hat{\mathbf{x}}^{k+1}$ obtained from Equation (26), we can further estimate $\bm{\beta}^{k+1}$. Rearrange the objective function as:
$\mathcal{L}\left(\hat{\mathbf{x}},\bm{\beta}^{k}\right)=$ $\frac{1}{2}\biggl(\bm{\beta}^{\mathrm{T}}\mathrm{diag}(\hat{\mathbf{x}}^{k+1})^{*}\mathbf{\mathcal{B}}^{\mathrm{H}}\mathbf{\mathcal{B}}\mathrm{diag}(\hat{\mathbf{x}}^{k+1})\bm{\beta}\biggr)$
$+\left((\hat{\mathbf{x}}^{k+1})^{\mathrm{H}}\mathbf{\mathcal{A}}^{\mathrm{H}}-(\mathbf{w}^{k})^{\mathrm{H}}\right)\mathbf{\mathcal{B}}\mathrm{diag}(\hat{\mathbf{x}}^{k+1})\bm{\beta}\biggr)+C^{\prime},$ (30)
where $C^{\prime}$ represents the sum of all terms that are not related to $\bm{\beta}$. Taking the first-order derivative of (30) with respect to $\bm{\beta}$ and setting it to zero, we have:
$=$ $\Re\left(\mathrm{diag}\left((\hat{\mathbf{x}}^{k+1})^{*}\right)\mathcal{B}^{\mathrm{H}}\left[\mathbf{w}^{k}-\mathcal{A}\hat{\mathbf{x}}^{k+1}\right]\right).$ (31)
Thus, we have the estimate of $\bm{\beta}^{k+1}$, given in Equation (27).
The proposed SBRI for the one-bit off-grid model is summarized in Algorithm 2.
@@
3:Initialize $\hat{\mathbf{x}}^{0}=\frac{\mathcal{A}^{\mathrm{H}}\overline{\mathbf{y}}}{\|\mathcal{A}^{\mathrm{H}}\overline{\mathbf{y}}\|_{2}}$, $\bm{\beta}^{0}=\bm{0}$
4:for $k=0$ to $T_{\max}-1$ do
5:  Compute $\mathbf{w}^{k}$ via (25): $\mathbf{w}^{k}\leftarrow[\Re(\overline{\mathbf{y}})\odot\Re(\mathbf{E}^{k}-\mathrm{I}^{\prime}(\mathbf{E}^{k}))]+j[\Im(\overline{\mathbf{y}})\odot\Im(\mathbf{E}^{k}-\mathrm{I}^{\prime}(\mathbf{E}^{k}))]$.
6:  Update $\hat{\mathbf{x}}^{k+1}$ via (26): $\hat{\mathbf{x}}^{k+1}\leftarrow\bigl(\mathbf{C}^{\mathrm{H}}(\bm{\beta}^{k})\mathbf{C}(\bm{\beta}^{k})+\gamma^{k}\bm{\Lambda}(\hat{\mathbf{x}}^{k})\bigr)^{-1}\mathbf{C}^{\mathrm{H}}(\bm{\beta}^{k})\mathbf{w}^{k}$
7:  $\bm{\beta}^{k+1}\leftarrow\text{gridUpdate}_{\text{SBRI}}(\hat{\mathbf{x}}^{k+1},\mathcal{A},\mathcal{B},\mathbf{w}^{k})$, $\text{gridUpdate}_{\text{SBRI}}(\cdot)$ is defined in (27).
8:  $\gamma^{k+1}\leftarrow\gamma^{0}\|\hat{\mathbf{x}}^{k+1}\|_{2}$
[…]
10:  end if
11:end for
12:return $\hat{\mathbf{x}}^{\star}\leftarrow\hat{\mathbf{x}}^{k+1}$, $\bm{\beta}^{\star}\leftarrow\bm{\beta}^{k+1}$
@@ III-C One-Bit DOA Estimation via Augmented Framework: SBRI-X
[…]
@@
[…]
Following the same approach as in Section III-A, taking the negative logarithm of Equation (38) yields the following non-convex objective function:
@@
$=$ $\sum_{i=1}^{N}\ln{\left(1+\mathrm{a}\exp\left(-\mathrm{b}\overline{y}_{i}\left[\mathbf{a}_{i}^{\mathrm{T}}\tilde{\mathbf{x}}+\epsilon_{i}\right]\right)\right)}\$
@@
III-C1 Majorizing Function for $\mathcal{L}_{ongrid}^{\diamond}\left(\tilde{\mathbf{x}},\bm{\epsilon}\right)$
@@
By substituting (42) into (39) and applying smooth approximation on the second term in (39), we obtain the majorizing function $\bar{\mathcal{L}}_{ongrid}^{\diamond}\left(\tilde{\mathbf{x}},\bm{\epsilon}\right)$ of $\mathcal{L}_{ongrid}^{\diamond}\left(\tilde{\mathbf{x}},\bm{\epsilon}\right)$:
@@
We have the following theorem regarding the convergence of $\tilde{\mathbf{x}}^{\star}$ in the SBRI-X framework.
Theorem 2.
Suppose that $\tilde{\mathbf{x}}^{\diamondsuit}$ is an $s$-sparse vector. Assume that $\mathbf{\mathcal{A}}$ obeys the RIP[57] of order $2s$ with constant $\delta_{2s}<1$ and a smoothing parameter $\eta>0$. Let $\{\tilde{\mathbf{x}}^{k}\}$ be the sequence generated by Algorithm 2 with $0<\alpha\leq 1$. Then $\{\tilde{\mathbf{x}}^{(k)}\}$ has at least one convergent subsequence. When $\eta>0$, the limit $\tilde{\mathbf{x}}^{\dagger}$ of any convergent subsequence is a limit point of the problem in Equation (43) and satisfies
$\left\lVert\tilde{\mathbf{x}}^{\dagger}-\tilde{\mathbf{x}}^{o}\right\rVert_{2}\;\leq\;C_{1}^{\prime}\sqrt{\frac{2(\mathrm{a}+1)}{\mathrm{a}\mathrm{b}^{2}}}\;+\;C_{2}^{\prime}\,\sigma_{s}\!\bigl(\tilde{\mathbf{x}}^{{\dagger}}\bigr)_{2},$ (45)
where $C_{1}^{\prime},C_{2}^{\prime}>0$ depend only on $\delta_{2s}$ and initial point $\hat{\mathbf{x}}^{0}$.
The proof of Theorem 2 is included in Appendix B.
@@ III-C2 Updating Formula
2:Output: $\tilde{\mathbf{x}}^{\star}$
3:Initialize $\tilde{\mathbf{x}}^{0}=\frac{\mathcal{A}^{H}\overline{\mathbf{y}}}{\|\mathcal{A}^{H}\overline{\mathbf{y}}\|_{2}}$
4:for $k=0$ to $T_{\max}-1$ do
5:  $\tilde{\mathbf{x}}^{k+1}\!\leftarrow\!(\mathcal{A}^{H}\mathcal{A}+\tfrac{\gamma^{k}(a+1)^{2}}{ab^{2}}\bm{\Lambda}(\tilde{\mathbf{x}}^{k}))\!^{-1}\!\mathcal{A}^{\mathrm{H}}[\mathcal{A}\tilde{\mathbf{x}}^{k}+g(\tilde{\mathbf{x}}^{k},\bm{\epsilon}^{k})]$
6:  $\bm{\epsilon}^{k+1}\!\leftarrow\!\bm{\epsilon}^{k}-\mathcal{A}(\tilde{\mathbf{x}}^{k+1}-\tilde{\mathbf{x}}^{k})+g(\tilde{\mathbf{x}}^{k},\bm{\epsilon}^{k})$
7:  $\gamma^{k+1}\leftarrow\gamma^{0}\|\tilde{\mathbf{x}}^{k+1}\|_{2}$
[…]
9:  end if
10:end for
11:return $\tilde{\mathbf{x}}^{\star}\leftarrow\tilde{\mathbf{x}}^{k+1}$
@@
$\tilde{\mathbf{x}}^{k+1}=$ $\left(\mathbf{C}^{\mathrm{H}}(\bm{\beta}^{k})\mathbf{C}(\bm{\beta}^{k})+\frac{\gamma^{k}(\mathrm{a}+1)^{2}}{\mathrm{a}\mathrm{b}^{2}}\bm{\Lambda}(\tilde{\mathbf{x}}^{k})\right)^{-1}$
@@
$g^{\dagger}(\tilde{\mathbf{x}}^{k},\bm{\epsilon}^{k},\bm{\beta}^{k})=\frac{(\mathrm{a}+1)^{2}\Re({\overline{\mathbf{y}}})}{\mathrm{b}\exp\left(\mathrm{b}\Re({\overline{\mathbf{y}}})\odot\Re\left(\mathbf{C}(\bm{\beta}^{k})\tilde{\mathbf{x}}^{k}+\bm{\epsilon}^{k}\right)\right)+\mathrm{a}\mathrm{b}}$
@@
2:Output: $\tilde{\mathbf{x}}^{\star}$, $\bm{\beta}^{\star}$
3:Initialize $\tilde{\mathbf{x}}^{0}=\frac{\mathcal{A}^{\mathrm{H}}\overline{\mathbf{y}}}{\|\mathcal{A}^{\mathrm{H}}\overline{\mathbf{y}}\|_{2}}$, $\bm{\beta}^{0}=\mathbf{0}$
4:for $k=0$ to $T_{\max}-1$ do
5:  $\tilde{\mathbf{x}}^{k+1}\!\leftarrow\!\left(\mathbf{C}^{\mathrm{H}}(\bm{\beta}^{k})\mathbf{C}(\bm{\beta}^{k})+\frac{\gamma^{k}(\mathrm{a}+1)^{2}}{\mathrm{a}\mathrm{b}^{2}}\bm{\Lambda}(\tilde{\mathbf{x}}^{k})\right)^{-1}$
[…]
7:  $\bm{\epsilon}^{k+1}\leftarrow\bm{\epsilon}^{k}-\mathbf{C}(\bm{\beta}^{k})\left(\tilde{\mathbf{x}}^{k+1}-\tilde{\mathbf{x}}^{k}\right)+g^{\dagger}(\tilde{\mathbf{x}}^{k},\bm{\epsilon}^{k},\bm{\beta}^{k})$
8:  $\bm{\beta}^{k+1}\!\leftarrow\!\text{gridUpdate}_{\text{SBRI-X}}(\tilde{\mathbf{x}}^{k+1},\bm{\epsilon}^{k+1},\tilde{\mathbf{x}}^{k},\bm{\epsilon}^{k},\mathcal{A},\mathcal{B})$, $\text{gridUpdate}_{\text{SBRI-X}}(\cdot)$ is defined in (53) and (54)
9:  $\gamma^{k+1}\leftarrow\gamma^{0}\|\tilde{\mathbf{x}}^{k+1}\|_{2}$
[…]
11:  end if
12:end for
13:return $\tilde{\mathbf{x}}^{\star}\leftarrow\hat{\mathbf{x}}^{k+1}$, $\bm{\beta}^{\star}\leftarrow\bm{\beta}^{k+1}$